# Fibred surfaces and their unitary rank

### v1

@@ 4.3 Slope inequalities for Harder-Narashiman sequence of lenght $\leq 2$
The result above gives a high bound for the case in which the Harder-Narasimhan sequence is of lenght $1$. It is not known wether this bound is sharp or not.
We are now going to give sharp bounds for the cases where the Harder-Narasimhan sequence of $\mathcal{E}=f_{*}\omega_{f}$ is of lenght $2$.
First we treat the case when $\mu_{l}=0$ (i.e. $u_{f}>0$).
@@
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $g\geq 2$, such that $\mathcal{A}$ is semistable and $\mu_{l}=0$ (so $u_{f}>1$). Then the following inequality holds:
@@
Moreover, assume that equality holds in (4.26); then:
    if $u_{f}\leq g-2$, then $F$ is hyperelliptic;
    if $u_{f}=g-1$, then $F$ is trigonal.
Assume that $u_{f}\leq c_{f}$. Then Then the following inequality holds:
@@
$K_{f}^{2}\geq\left(2g-1+\frac{2g-2-c_{f}}{g-c_{f}-1}(g-u_{f}-1)-\frac{g-c_{f}-1}{g(2u_{f}-2c_{f}+1)+u_{f}(g-c_{f}-2)-2}\right)\frac{\chi_{f}}{(g-u_{f})}.$ (4.28)
@@
By the assumptions the Harder-Narasimhan sequence of the Hodge bundle $\mathcal{E}$ has lenght 2 and is of this form:
@@
We use Theorem 4.12, observing that under our assumptions $r_{1}=g-u_{f}$, and (4.13) becomes
@@
The case where equality holds follows again from Theorem 4.12. As for the other two inequalities, they follow directly from the result of Theorem 4.16. ∎
@@
Observe that using Xiao’s method with the assumptions of Theorem 4.25 one obtains in this case (as explained for instance in [7, Remark 3.3]) the worse inequality:
@@
This very same equality is reproved in the article [17] in the proof of Lemma 3.3.2.
Although we will not present any consequence of this result, let us complete the picture by giving an inequality for the case when the Harder-Narashiman sequence has lenght 2 and $u_{f}=0$.
@@
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $g\geq 2$, such that $l=2$, and $\mu_{l}>0$ (so in particular $u_{f}=0$). Then the following inequality holds:
$K_{f}^{2}\geq\frac{(4(g-1)[(2g-2r_{1}+1)(g+r_{1}-2)+g-r_{1}}{2g^{3}+(2r_{1}-2)g^{2}+(-6{r_{1}}^{2}+9r_{1}-2)g+2{r_{1}}^{2}-3{r_{1}}^{2}}\chi_{f}.$ (4.31)
@@
We have that $r_{2}=g$, and (3.1) $\chi_{f}=\mu_{1}r_{1}+\mu_{2}(r_{2}-r_{1})$. By Lemma 3.16 and Remark 3.22 we have
$K_{f}^{2}\geq\left(\frac{(\alpha_{i}+1)(4g-4-a_{i})+a_{i}}{\alpha_{i}+1}\right)\mu_{i},$
for $i=1,2$.
For $i=1$ we have that $\alpha_{1}\leq a_{1}\leq 2g-2r_{1}$, so
$K_{f}^{2}\geq\frac{(2g-2r_{1}+1)(2g+2r_{1}-4)+2g-2r_{1}}{2g-2r_{1}+1}\mu_{1}.$
For $i=2$ we have $a_{2}=\alpha_{2}=0$, so we obtain
$K_{f}^{2}\geq 4(g-1)\mu_{2}$
So, putting these two inequalities together, we obtain
$\chi_{f}=\mu_{1}r_{1}+\mu_{2}(g-r_{1})\leq\left(\frac{2g-2r_{1}+1}{(2g-2r_{1}+1)(2g+2r_{1}-4)+2g-2r_{1}}r_{1}+\frac{g-r_{1}}{4(g-1)}\right)K_{f}^{2}$
and a straightforward computation gives the desired inequality. ∎
Remark 4.32.
The coefficient of $\chi_{f}$ is an increasing function of $r_{1}$, in the worst case $r_{1}=1$ we have
$K_{f}^{2}\geq\frac{8g(g-1)^{2}}{2g^{3}+g-1}\chi_{f},$
which is worse than the slope inequality. But for increasing values of $r_{1}$ the function gets better. For instance, in the best case $r_{1}=g-1$ we have
$K_{f}^{2}\geq\left(\frac{3(g-1)}{2}+\frac{1}{4(g-1)}\right)\chi_{f},$
which of course is extremely high.
5 A discussion on the extremal cases
We now list some results, due to Serrano, Pirola, Zucconi and Beorchia, about non locally trivial fibred surfaces with maximal unitary rank $u_{f}=g-1$
First of all, in [47, Theorem 3.6 and Remark 3.7] Pirola proved that in this case up to a finite base change $g-1=u_{f}=q_{f}$ (see Remark 2.23 to see that this is not obvious).
Proposition 5.1.
Let $f\colon S\to B$ a relatively minimal non trivial fibred surface of genus $g\geq 4$ such that $u_{f}=g-1$. Then the general fibre $F$ is trigonal;
Proof.
By Pirola’s result we can assume that the relative irregularity is maximal. Cai proved in [15] that for non locally trivial hyperelliptic fibrations Xiao’s bound (2.18) $q_{f}\leq(g+1)/2$ holds, so for $g\geq 3$ no hyperelliptic fibration can have $q_{f}=g-1$. In case $g=2$ of course $F$ is hyperelliptic. The case $g=3$ has been investigated in [44]. On the other hand, by Barja-Naranjo-Gonzalez [6] (later generalized to $u_{f}$ in [30]), we have that $q_{f}\leq g-c_{f}$ where $c_{f}=\mbox{{Cliff}}(F)$ is the Clifford index of the general fibre. Moreover Favale-Naranjo-Pirola [24] proved that if the general fibre is a degree $d$ plane curve (thus of gonality $d-1$ and Clifford index $d-3$ [3]), then $q_{f}\leq g-c_{f}-1$. So, under the assumption of the proposition, $\mbox{{Cliff}}(F)\leq 1$ and $F$ is not a plane quintic, nor it is hyperelliptic (i.e. with $\mbox{{Cliff}}(F)>0$). Hence we conclude that $F$ necessarily is trigonal. Observe that for this last part we could have used directly the unitary rank using the results of the author with Gonzàles-Alonso and Torelli in [30]. ∎
@@ 5 A discussion on the extremal cases
Let $f\colon S\to B$ a relatively minimal non trivial fibred surface of genus $g\geq 4$ such that $u_{f}=g-1$. Then the following hold:
@@
(1) This is due to Serrano: indeed, in [49] he proves that Xiao’s bound (2.18) $q_{f}\leq(g+1)/2$ holds for isotrivial fibrations. So we can exclude the isotrivial fibrations unless $g\leq 3$. In in [44] are indeed constructed also isotrivial fibrations of genus $2$ and $3$ with irregularity respectively $1$ and $2$.
(2) This proof is due to Pietro Pirola in a private communication. Let us consider a smooth fibre of $f$, $F_{t}$, for $t\in B^{0}$ (the smooth locus of $f$). We have the following diagram
@@
Now, the morphism ${j_{t}}_{*}$ goes in a family of translates of a fixed variety $A\subset\mbox{{Alb}}(S)$ of dimension $q_{f}=u_{f}=g-1$. Let us call $g_{t}\colon J(F_{t})\to A$ this morphism, fixing $A$. Then if we pass to the dual varieties and morphisms we have $(g_{t})^{*}\colon\widehat{A}\hookrightarrow\mbox{{Pic}}^{0}(F_{t})$. Let us consider the quotient abelian variety, which is an elliptic curve, and denote it $E_{t}$. We have the morphisms
$F_{t}\hookrightarrow\mbox{{Pic}}^{0}(F_{t})\longrightarrow E_{t},$
which is necessarily surjective because the image of $F_{t}$ generates $\mbox{{Pic}}^{0}(F_{t})$ as an abelian variety. So, this is the desired morphism of the general fibre to a family of elliptic curves.
Now, the family $E_{t}$ has varying moduli: if this was not the case than all $\mbox{{Pic}}^{0}(F_{t})$ would be a fixed abelian variety and so the entire family would be trivial up to a base change, hence isotrivial, but this is not the case for (2).
@@
${S^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}}$$\scriptstyle{\beta}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${W\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B}$
Now the fibration of genus one contains at least one fibre with non compact Jacobian. Thus $f$ contains at least a fibre with the same property, and this implies that it is not isotrivial.
The last conclusion could be derived also from the fact that a non-isotrivial family of curves of genus $1$ contains at least a vertical rational curve: see [2]. So necessarily $\alpha$ contains fibres whose support has more than one component. This implies that also $f^{\prime}$ necessarily has the same feature, and so also $f$ necessarily contains singular fibres. ∎
@@ 6 An upper bound on the unitary rank
We now extend Konno’s bound [34, Lemma 2.7] on $q_{f}$ to $u_{f}$, and improve it obtaining a strict inequality.
@@
The non-sharp inequality is immediate by by combining inequalities (4.3) and (2.11). Suppose that the equality equality holds in (6.2), then and equality in (4.3) holds, so $f$ is necessarily of maximal unitary rank $u_{f}=g-1$ and equality in (2.11) holds, so $f$ is necessarily Kodaira. We have seen that this is impossible in Theorem 5.2, thus the inequality is always strict. ∎
@@
Let $f\colon S\to B$ be a relatively minimal non-isotrivial fibration of genus $g\geq 2$. Then Let us suppose that $u_{f}$ is maximal, i.e. $u_{f}=g-1$. Then $g\leq 6$
@@
Inequality (4.3) in the first case becomes
@@
If we combine it with (2.11), we obtain $4g(g-1)\leq 3(2g-1)$, so $g^{2}-7g+3\leq 0$, which gives that $g\leq 6$. ∎
We can also give a bound in the case $u_{f}=g-2$. By using equation (6.2) in this case we obtain the equation
@@
which implies $g\leq 12$, but we now see that we can rule out case $g=12$ as follows.
@@
and so, combining again with (2.11) we get $g\leq 12$. But for $g=12$ we obtain that equality is reached in (4.3), so $u_{f}$ should be $g-1$, a contradiction. ∎
Remark 6.5.
One could also use the results of Section 5 to exclude the case $g=12$.
I the last section 8 we give many conditions on the case when $g=6$ and $u_{f}=g-1=5$.
@@ 7 An application to the Coleman-Oort conjecture
We will use the setting and the notations of the papers [40] and [17]. Let us consider $\mathcal{M}_{g}=\mathcal{M}_{g,\ell}$ the moduli space of of smooth projective curves of genus $g\geq 2$ with a full level $\ell-$structure. and $\mathcal{A}_{g}=\mathcal{A}_{g,\ell}$ the moduli space of $g$-dimensional principally polarized abelian varieties with a full level $\ell$-structure. The level $\ell\geq 3$ assures that these are moduli schemes representing the corresponding moduli functor. Let
@@
The moduli space $\mathcal{A}_{g}$ is isomorphic to a Shimura variety. In $\mathcal{A}_{g}$ there are Shimura subvarieties. The famous Coleman-Oort conjecture predicts that for large $g$ there do not exist any Shimura subvariety generically contained in the Torelli locus.
@@
Here, using the results of the previous sections, we can give an improvement of a result of Chen, Lu and Zuo in [17, Thm 1.1.2], giving a condition on the rank of the $(-1,0)$ part of the maximal unitary Higgs subbundle $\mathcal{U}_{C}$ of a curve $C\subset\mathcal{A}_{g}$ that implies that it is not generically contained in the Torelli locus.
@@
The proof is completely analogous to the one of [17, Thm 1.12]. The key point is that if $C$ is contained generically in the Torelli locus, then we can consider the following diagram
@@
where $B$ is the normalization of the pullback $(j^{\circ})^{-1}C$ in $\mathcal{M}_{g}$. The morphism $B\to\mathcal{M}_{g}$ gives a holomorphic fibration $f\colon S\to B$ induced by the fine moduli structure. We can perform the semistable reduction to a compactification obtaining a non-isotrivial fibred surface $\overline{f}\colon\overline{S}\to\overline{B}$, and a morphism $\overline{B}\to\overline{\mathcal{M}}_{g}$ that extends $B\to\mathcal{M}_{g}$.
Now, the Hodge bundle $\overline{f}_{*}\omega_{\overline{f}}$ is the $(-1,0)-$part of the Higgs bundle on $\overline{B}$ associated to the relative semistable Jacobian family (see [38, Sec. 3]). The pullback via $j_{\overline{B}}$ of the $(-1,0)$ part of the decomposition $\mathcal{E}_{\overline{B}}=\mathcal{F}_{\overline{B}}\oplus\mathcal{U}_{\overline{B}}$ is precisely the second Fujita decomposition (2.7) of ${\overline{f}}_{*}\omega_{\overline{f}}$, so
@@
Now the first statement $(i)$ directly follows from Theorem 6.1. For the second statement $(ii)$ one needs to use a result due to Viehweg and Zuo ([50], see also [17, Thm. 1.2.1]), that tells us in particular that for the case when $C$ is Shimura, then $\mathcal{F}_{\overline{B}}^{-1,0}$ is semistable. So, in this case we are in the assumptions of Theorem 4.25, and so we have inequality (4.26), that combined with (2.11) gives:
@@
If equality holds, then in particular $f$ is Kodaira, and equality holds in (4.26), so by Theorem 4.25 we have that:
  • either $u_{\overline{f}}=g-1$, but in this case we already proved in Theorem 5.2 that the fibration is not Kodaira;
  • or $u_{\overline{f}}\leq g-2$ and the fibration is hyperelliptic. In this case we know (see (U4) of [30]) that $u_{\overline{f}}\leq(g+1)/2$, so we get that this is impossible for $g\leq 4$. But also for $g=2,3$ the equality in (7.4) cannot hold, because the discriminant is not a square number.
@@
The bounds found by Chen, Lu and Zuo in [17, Thm 1.12] are respectively:

### v2

@@ 4.3 Slope inequalities for Harder-Narasimhan filtration of length $\leq 2$
The result above gives a high bound for the case in which the Harder-Narasimhan filtration is of length $1$. It is not known whether this bound is sharp or not.
We now give a new bound for the slope under the assumption that Harder-Narasimhan filtration of $\mathcal{E}=f_{*}\omega_{f}$ has length $2$ and that $\mu_{l}=0$ (i.e., $u_{f}>0$).
@@
Let $f\colon S\to B$ be a relatively minimal locally non-trivial fibred surface of genus $g\geq 2$, such that $\mathcal{A}$ is semistable and $\mu_{l}=0$. Then the following inequality holds:
@@
Moreover, assume that the equality holds in (4.27). Then necessarily $u_{f}=g-1$ and $F$ is trigonal.
Assume that $u_{f}\leq c_{f}$. Then the following inequality holds:
@@
$K_{f}^{2}\geq\left(2g-1+\frac{2g-2-c_{f}}{g-c_{f}-1}(g-u_{f}-1)-\frac{g-c_{f}-1}{2gu_{f}-gc_{f}-2u_{f}-c_{f}u_{f}-1}\right)\frac{\chi_{f}}{(g-u_{f})}.$ (4.29)
@@
From the assumptions we have that the Harder-Narasimhan filtration of the Hodge bundle $\mathcal{E}$ has length 2:
@@
We use Theorem 4.13: the inequality (4.14) becomes
@@
The case where the equality holds follows again from Theorem 4.13 (recalling that in this case $r_{1}<g$). As for the other two inequalities, they follow directly from Theorem 4.18. ∎
@@
As explained for example in [4, Remark 3.3], using Xiao’s method with the assumptions of Theorem 4.26 one obtains the worse inequality:
@@
This very same inequality is reproved in the article [10] in the proof of Lemma 3.3.2.
5 A discussion on the extremal cases
We now list some results, due to Serrano, Pirola, Zucconi and Beorchia, about locally non-trivial fibred surfaces with maximal unitary rank $u_{f}=g-1$
First of all, in [35, Theorem 3.6 and Remark 3.7] Pirola proved that in this case up to an étale base change $g-1=u_{f}=q_{f}$ (see Remark 2.23 to see that this is not obvious).
@@ 5 A discussion on the extremal cases
Let $f\colon S\to B$ a relatively minimal non-trivial fibred surface of genus $g\geq 4$ such that $u_{f}=g-1$. Then the general fibre $F$ is trigonal;
@@
By Pirola’s result we can assume that the relative irregularity is maximal. Cai proved in [8] that for locally non-trivial hyperelliptic fibrations Xiao’s bound (2.18) $q_{f}\leq(g+1)/2$ holds, so for $g\geq 3$ no hyperelliptic fibration can have $q_{f}=g-1$. In the case where $g=2$ of course $F$ is hyperelliptic. The case $g=3$ has been investigated in [32]. On the other hand, by Barja-Naranjo-Gonzalez [3], we have that $q_{f}\leq g-c_{f}$. Moreover, Favale-Naranjo-Pirola [14] proved that if the general fibre is a degree $d$ plane curve (thus of gonality $d-1$ and Clifford index $d-3$), then $q_{f}\leq g-c_{f}-1$. Therefore, under the assumption of the proposition, $\mbox{{Cliff}}(F)\leq 1$ and $F$ is not a plane quintic, nor it is hyperelliptic (i.e., with $\mbox{{Cliff}}(F)>0$). Hence we conclude that $F$ necessarily is trigonal. ∎
@@
Let $f\colon S\to B$ be a relatively minimal non-trivial fibred surface of genus $g\geq 4$ such that $u_{f}=g-1$. Then the following hold:
@@
By Pirola’s result we can assume that after an étale base change, $q_{f}=u_{f}=g-1$. Étale base changes do not modify the slope.
(1) This is due to Serrano: indeed, in [37] he proved that Xiao’s bound (2.18) $q_{f}\leq(g+1)/2$ holds for isotrivial fibrations. So we can exclude the isotrivial fibrations unless $g\leq 3$. In [32] isotrivial fibrations of genus $2$ and $3$ with irregularity respectively $1$ and $2$ are constructed.
(2) This proof is due to Pietro Pirola in a private communication. Let $B^{0}$ be the smooth locus of $f$ (i.e., the locus consisting of all $t\in B$ such that the fibre $F_{t}$ is smooth). For any $t\in B^{0}$, let $j_{t}\colon F_{t}\hookrightarrow S$ be the inclusion. We have the following diagram
@@
Now, the morphism ${j_{t}}_{*}$ goes in a family of translates of a fixed abelian sub-variety $A\subset\mbox{{Alb}}(S)$ of dimension $q_{f}=u_{f}=g-1$. Let us consider the induced surjective homomorphism $g_{t}\colon J(F_{t})\to A$, and let be $(\ker(g_{t})_{0})$ be the abelian subvarieties of $J(F_{t})$ that are the connected component of $\ker(g_{t})$ containing the origin, for any $t\in B^{0}$. Note that $(\ker(g_{t})_{0})$ is an elliptic curve for any $t\in B^{0}$. Passing to the dual varieties and homomorphisms we have surjective homomorphisms
$\mbox{{Pic}}^{0}(F_{t})\to(\ker(g_{t})_{0})^{\vee}=:E_{t}$
for any $t\in B^{0}$. We thus have the morphisms
$F_{t}\hookrightarrow\mbox{{Pic}}^{0}(F_{t})\longrightarrow E_{t}.$
The composition $F_{t}\to E_{t}$ necessarily is surjective because the image of $F_{t}$ generates $\mbox{{Pic}}^{0}(F_{t})$ as an abelian variety. Therefore, this is the desired morphism of the general fibre to a family of elliptic curves.
Now, the family $E_{t}$ has varying moduli: if this were not the case then all $\mbox{{Pic}}^{0}(F_{t})$ would be a fixed abelian variety. This would imply that the entire family over $B_{0}$ would be trivial up to a base change, a contradiction by (1).
@@
[…]
Now the fibration of genus one contains at least one fibre with non compact Jacobian. Thus $f$ contains at least a fibre with the same property, and this implies that it is not isotrivial. ∎
@@ 6 An upper bound on the unitary rank
We now extend Konno’s bound [21, Lemma 2.7] on $q_{f}$ to $u_{f}$, and improve it to a strict inequality.
@@
The non-sharp inequality is immediate by combining the inequalities (4.3) and (2.12). Suppose now that the equality holds in (6.2). Then the equality in (4.3) holds, so $f$ is necessarily of maximal unitary rank $u_{f}=g-1$, and the equality in (2.12) holds; so $f$ is necessarily Kodaira. We have seen in Theorem 5.2 that this case cannot happen. ∎
@@
Let $f\colon S\to B$ be a relatively minimal non-isotrivial fibration of genus $g\geq 2$. Suppose that $u_{f}$ is maximal, i.e., $u_{f}=g-1$. Then $g\leq 6$.
@@
The inequality (4.3) in the first case becomes
@@
If we combine it with the bound (2.12), we obtain $4g(g-1)\leq 3(2g-1)$, so $g^{2}-7g+3\leq 0$, which gives that $g\leq 6$. ∎
We can also give a bound in the case $u_{f}=g-2$. By using the inequality (6.2) in this case we obtain the equation
@@
which implies $g\leq 12$. We can rule out the case where $g=12$ as follows.
@@
and so, combining again with (2.12) we get $g\leq 12$. But for $g=12$ we obtain that the equality is reached in (4.3) of Theorem 4.1, so $u_{f}$ should be $g-1$, a contradiction. ∎
I the paper [39] we give many conditions on the case where $g=6$ and $u_{f}=g-1=5$.
@@ 7 An application to the Coleman-Oort conjecture
We will use the setting and the notations of the papers [26] and [10]. Let us consider $\mathcal{M}_{g}=\mathcal{M}_{g,\ell}$ the moduli space of smooth projective curves of genus $g\geq 2$ with a full level $\ell-$structure. and $\mathcal{A}_{g}=\mathcal{A}_{g,\ell}$ the moduli space of $g$-dimensional principally polarized abelian varieties with a full level $\ell-$structure. The level $\ell\geq 3$ assures that these are moduli schemes representing the corresponding moduli functor. Let
@@
The moduli space $\mathcal{A}_{g}$ is isomorphic to a Shimura variety. In $\mathcal{A}_{g}$ there are Shimura subvarieties. The famous Coleman-Oort conjecture predicts that for large $g$ there does not exist any positive-dimensional Shimura subvariety generically contained in the Torelli locus.
@@
Here, using the results of the previous sections, we can improve a result of Chen, Lu and Zuo in [10, Thm 1.1.2]. We give a condition on the rank of the $(-1,0)$ part of the maximal unitary Higgs subbundle $\mathcal{U}_{C}$ of a curve $C\subset\mathcal{A}_{g}$ that implies that it is not generically contained in the Torelli locus.
@@
The proof is completely analogous to the one of [10, Thm 1.12]. The key point is that if $C$ is contained generically in the Torelli locus, then we can consider the following diagram
@@
where $B$ is the normalization of the pullback $(j^{\circ})^{-1}C$ in $\mathcal{M}_{g}$. The morphism $B\to\mathcal{M}_{g}$ gives a fibration $f\colon S\to B$ induced by the fine moduli structure. We can perform the semistable reduction to a compactification of the family obtaining a non-isotrivial fibred surface $\overline{f}\colon\overline{S}\to\overline{B}$, and a morphism $\overline{B}\to\overline{\mathcal{M}}_{g}$ extending $B\to\mathcal{M}_{g}$.
Now, the Hodge bundle $\overline{f}_{*}\omega_{\overline{f}}$ is the $(-1,0)-$part of the Higgs bundle on $\overline{B}$ associated to the relative semistable Jacobian family (see [24, Sec. 3]). The pullback via $j_{\overline{B}}$ of the $(-1,0)$ part of the decomposition $\mathcal{E}_{\overline{B}}=\mathcal{F}_{\overline{B}}\oplus\mathcal{U}_{\overline{B}}$ is precisely the second Fujita decomposition (2.7) of ${\overline{f}}_{*}\omega_{\overline{f}}$, so
@@
The statement $(i)$ directly follows from Theorem 6.1.
In order to prove $(ii)$, we need a result due to Viehweg and Zuo ([40], see also [10, Thm. 1.2.1]): if $C$ is Shimura, then $\mathcal{F}_{\overline{B}}^{-1,0}$ is semistable. We thus are in the assumptions of Theorem 4.26. The inequality (4.27) combined with (2.12) gives:
@@
If the equality holds, then in particular $f$ is Kodaira, and the equality holds in (4.27), so by Theorem 4.26 we have that: $u_{\overline{f}}=g-1$. In this case we already proved in Theorem 5.2 that the fibration is not Kodaira. ∎
@@
The bounds found by Chen, Lu and Zuo in [10, Thm 1.12] are respectively: